· The voule of a glass vessel is 1000 cc at 20 °C.What voulme of mercury should be poured into it at this temperature so that the volumeof the remaining space does not change with temperature? Coefficients of cubical expansion of mercury and glass are 1.8x10^-4/°C and 9x10^-6/°C respectively.



·         The voule of a glass vessel is 1000 cc at 20



  • °C.What voulme of mercury should be poured into it at this temperature so that the volumeof the remaining space does not change with temperature? Coefficients of cubical expansion of mercury and glass are 1.8x10^-4/°C and 9x10^-6/°C respectively.



Grade:

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Cathy,

Ans:- Let the vol of the vessel is V and that of the mercury is xV where x is a fractio whose value we have to determine.

Now if the expansion of both of them are same then there will be no change in the apparent volume of the empty portion.

let the temp diff is t . So, using the same condition we get,

ΔV(glass)=ΔV(mercury)

or V×9×10^-6×t=(xV)×1.8×10^-4

x=1/20

Hence the vol of murcury should be taken=1000×1/20=50cc(ans)

 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Cathy !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur